OpenStudy (dominickn609):

i need help with evaluating expressions and i have no idea how to can some one please help

OpenStudy (mathmate):

Please clarify what kind of expressions... algebraic, numeric, trigonometric... For numeric expressions, use PEDMAS (or PEMDAS), see, for example: http://www.purplemath.com/modules/orderops.htm

OpenStudy (dominickn609):

algebraic i think

OpenStudy (dominickn609):

\[(2 \times 3)^2 \times -8 - 4(4+10)^2 + 8\]

OpenStudy (mathmate):

This is numeric, you would use PEMDAS. The order of evaluation is very important, PEMDAS stands for P=parentheses, anything inside brackets are evaluated first E=exponentiation, like squares, cubes, etc. are evaluated next. MD = multiplications and divisions, next in order. If multiplications & divisions follows one another, do from left to right. AS = addition and subtraction, last in order. If they follow one another, do from left to right. Again, http://www.purplemath.com/modules/orderops.htm explains clearly with examples. Example: \((3 + 4)^2 \times (2) - 4(1+2)^2 + 3\) evaluate expressions inside brackets: \((7)^2\times 2 -4(3)^2+3\) next, do exponentiation \(49\times 2-4\times 9+3\) next do multiplications (and divisions, if any) 98-36+3 next do additions and subtractions, from left to right 62+3 =65 is the final answer.